OpenStudy (anonymous):

Martin drew a pair of perpendicular lines and a pair of parallel lines. Which of these statements best compares the pairs of perpendicular and parallel lines?

OpenStudy (anonymous):

from equation 1.... \[x=3+y\] putting in 2nd equ \[2(3+y)+2y=2\] \[6+2y+2y=2\] \[6+4y=2\]\[4y=2-6\] \[4y=-4\] \[y=-1\] for x ut y=-1 in equ 1 \[x-y=3\] \[x-(-1)=3\] \[x+1=3\]\[x=2\]